OpenStudy (anonymous):

Write the equation of each line in slope-intercept form, y=mx+b. Line passing through f(1)= 1 and f(-1)= 2

OpenStudy (anonymous):

You can make the equations: \[1=m*1+b\]\[2=m*(-1)+b\]

OpenStudy (triciaal):